# Extraction of ground-state nuclear deformations from ultra-relativistic heavy-ion collisions: Nuclear structure physics context

### v1

@@ VI Conclusions
First, the modeling of the initial QGP state formed following the heavy-ion collision involves the description of the relative position identification of the individual nucleons. To this end, the information contained in the one-body nucleonic density, as presented in the majority of papers dealing with nuclear imaging from UHIC, is insufficient. The proper tool for this imaging of the nucleon’s position distributions is the two-body density (to map axial distributions) or the three-body one (to map triaxial shapes). The schematics depicting the deformed orientations of nuclei are misleading, as even-even nuclei, such as 238U or 20Ne, have isotropic ground-state wave functions in the laboratory system.
@@
Fourth, the existing precise information on nuclear shapes should be utilized by nuclear UHIC imaging practitioners to validate analysis techniques and control model uncertainties and discrepancies. Currently, a comprehensive and combined uncertainty quantification analysis has not been carried out.
In summary, we find that the impact of nuclear UHIC imaging on low-energy nuclear structure is still limited. Nuclear shape parameters for stable or long-lived nuclei are well understood, having been studied extensively over many years using a wide range of techniques, both direct and indirect. Given the significant uncertainties associated with the UHIC imaging process, which relies on phenomenological models, the accuracy of current systematic nuclear structure measurements and calculations appears sufficient to inform the modeling of the QGP initial state.
@@ Acknowledgements.
This material is based upon work supported by the U.S. Department of Energy under Award Numbers DE-SC0013365 and DE-SC0023633 (Office of Science, Office of Nuclear Physics), DE-SC0023175 (Office of Science, NUCLEI SciDAC-5 collaboration), DE-FG02-97ER41041, DE-FG02-97ER41033, and the National Science Foundation under award number 2004601 (CSSI program, BAND collaboration). This work was also partially supported by the STFC Grant Nos. ST/P003885/1 and ST/V001035/1.References
@@
- The STAR Collaboration [2025] The STAR Collaboration, Imaging nuclear shape through anisotropic and radial flow in high-energy heavy-ion collisions (2025), arXiv:2506.17785 [nucl-ex] .
@@
- Jia [2022] J. Jia, Shape of atomic nuclei in heavy ion collisions, Phys. Rev. C 105, 014905 (2022).
@@
- Giacalone et al. [2024] G. Giacalone, B. Bally, G. Nijs, S. Shen, T. Duguet, J.-P. Ebran, S. Elhatisari, M. Frosini, T. A. Lähde, D. Lee, B.-N. Lu, Y.-Z. Ma, U.-G. Meißner, J. Noronha-Hostler, C. Plumberg, T. R. Rodríguez, R. Roth, W. van der Schee, and V. Somà, The unexpected uses of a bowling pin: exploiting 20Ne isotopes for precision characterizations of collectivity in small systems (2024), arXiv:2402.05995 [nucl-th] .
- Giacalone et al. [2025] G. Giacalone, W. Zhao, B. Bally, S. Shen, T. Duguet, J.-P. Ebran, S. Elhatisari, M. Frosini, T. A. Lähde, D. Lee, B.-N. Lu, Y.-Z. Ma, U.-G. Meißner, G. Nijs, J. Noronha-Hostler, C. Plumberg, T. R. Rodríguez, R. Roth, W. van der Schee, B. Schenke, C. Shen, and V. Somà, Anisotropic flow in fixed-target Ne collisions as a probe of quark-gluon plasma, Phys. Rev. Lett. 134, 082301 (2025).
- Li et al. [2025] Y. Li, X. Zhang, G. Giacalone, and J. Yao, Benchmarking nuclear matrix elements of decay with high-energy nuclear collisions (2025), arXiv:2502.08027 [nucl-th] .
@@
- Zhou and Li [2025] Q. Zhou and Z. Li, Masses and Quadrupole Deformations of Even-Z Nuclei Within a Triaxial Relativistic Hartree–Bogoliubov Model, Particles 8, 10.3390/particles8020057 (2025).
@@
- Vernon et al. [2022] A. R. Vernon, R. F. Garcia Ruiz, T. Miyagi, C. L. Binnersley, J. Billowes, M. L. Bissell, J. Bonnard, T. E. Cocolios, J. Dobaczewski, G. J. Farooq-Smith, K. T. Flanagan, G. Georgiev, W. Gins, R. P. de Groote, R. Heinke, J. D. Holt, J. Hustings, Á. Koszorús, D. Leimbach, K. M. Lynch, G. Neyens, S. R. Stroberg, S. G. Wilkins, X. F. Yang, and D. T. Yordanov, Nuclear moments of indium isotopes reveal abrupt change at magic number 82, Nature 607, 260 (2022).
- Sun et al. [2025] Z. H. Sun, A. Ekström, C. Forssén, G. Hagen, G. R. Jansen, and T. Papenbrock, Multiscale physics of atomic nuclei from first principles, Phys. Rev. X 15, 011028 (2025).

### v3

@@ VI Conclusions
First, the modeling of the initial QGP state formed following the heavy-ion collision involves the description of the relative position identification of the individual nucleons. To this end, the information contained in the one-body nucleonic density, as presented in the majority of papers dealing with nuclear imaging from UHIC, is insufficient. The proper tool for this imaging of the nucleon’s position distributions is the two-body density (to map axial distributions) or the three-body one (to map triaxial shapes). The schematics depicting the deformed orientations of nuclei are misleading, as even-even nuclei, such as 238U or 20Ne, have isotropic ground-state wave functions in the laboratory system. In this context, the recently proposed approaches in Refs. [33, 34, 35, 19, 28], based on the many-body wave function, hold promise.
@@
Fourth, the existing precise information on nuclear shapes should be utilized by nuclear UHIC imaging practitioners to validate analysis techniques and control model uncertainties and discrepancies. Currently, a comprehensive and combined uncertainty quantification analysis of UHIC data has not been carried out because of the lack of a consistent methodology for propagating uncertainties from nuclear models through to final-state observables, and with propagating uncertainties backwards in the pipeline.
In summary, we find that the current use of UHIC to image low-energy nuclear structure – namely, nuclear deformations – is prone to many flaws of interpretation and precision. Nuclear one-body electric moments for stable or long-lived nuclei are well understood and have been studied extensively over many years using a wide range of techniques, both direct and indirect. This is still a very active area of research, as many detailed challenges remain in low-energy experiment and theory. The UHIC measurements do not directly contribute to this field of study because they provide many-body correlations in the laboratory frame. As proposed in this work, there is potential for using the measured many-body correlations by the UHIC nuclear imaging to examine multipole collectivity from an entirely new perspective. This could open up a fascinating new research avenue.
The modeling dependence, large uncertainties, and limited isotopic reach of UHIC-based studies pose significant challenges and hinder the method’s ability to meaningfully add to the existing corpus of data on nuclear shapes. On the other hand, the existing systematic nuclear structure measurements and calculations provide strong constraints for the modeling of the initial state of UHIC and a firm baseline for future studies of QGP formation and flow. With that being said, a compelling feature of UHIC analyses is that they explicitly probe many-body correlations in the nucleus. Focusing on this aspect, rather than shapes, represents a strength of the approach worth playing to.
@@ Acknowledgements.
This material is based upon work supported by the U.S. Department of Energy under Award Numbers DE-SC0013365 and DE-SC0023633 (Office of Science, Office of Nuclear Physics), DE-SC0023175 (Office of Science, NUCLEI SciDAC-5 collaboration), DE-FG02-97ER41041, DE-FG02-97ER41033, and the National Science Foundation under award number 2004601 (CSSI program, BAND collaboration). This work was also partially supported by the STFC Grant Nos. ST/V001035/1 and ST/Y000285/1.References
- Alver et al. [2008] B. Alver et al., Importance of correlations and fluctuations on the initial source eccentricity in high-energy nucleus-nucleus collisions, Phys. Rev. C 77, 014906 (2008).
@@
- Adamczyk et al. [2015] L. Adamczyk et al. (STAR Collaboration), Azimuthal anisotropy in and collisions at RHIC, Phys. Rev. Lett. 115, 222301 (2015).
@@
- Jia [2022a] J. Jia, Shape of atomic nuclei in heavy ion collisions, Phys. Rev. C 105, 014905 (2022a).
- Jia [2022b] J. Jia, Probing triaxial deformation of atomic nuclei in high-energy heavy ion collisions, Phys. Rev. C 105, 044905 (2022b).
@@
- STAR Collaboration [2025] STAR Collaboration, Imaging nuclear shape through anisotropic and radial flow in high-energy heavy-ion collisions, Rep. Prog. Phys. 10.1088/1361-6633/ae0fc3 (2025).
- Giacalone et al. [2025a] G. Giacalone, B. Bally, G. Nijs, S. Shen, T. Duguet, J.-P. Ebran, S. Elhatisari, M. Frosini, T. A. Lähde, D. Lee, B.-N. Lu, Y.-Z. Ma, U.-G. Meißner, J. Noronha-Hostler, C. Plumberg, T. R. Rodríguez, R. Roth, W. van der Schee, and V. Somà, Exploiting isotopes for precision characterizations of collectivity in small systems, Phys. Rev. Lett. 135, 012302 (2025a).
- Giacalone et al. [2025b] G. Giacalone, W. Zhao, B. Bally, S. Shen, T. Duguet, J.-P. Ebran, S. Elhatisari, M. Frosini, T. A. Lähde, D. Lee, B.-N. Lu, Y.-Z. Ma, U.-G. Meißner, G. Nijs, J. Noronha-Hostler, C. Plumberg, T. R. Rodríguez, R. Roth, W. van der Schee, B. Schenke, C. Shen, and V. Somà, Anisotropic flow in fixed-target Ne collisions as a probe of quark-gluon plasma, Phys. Rev. Lett. 134, 082301 (2025b).
- Li et al. [2025] Y. Li, X. Zhang, G. Giacalone, and J. Yao, Benchmarking nuclear matrix elements of decay with high-energy nuclear collisions, Phys. Rev. Lett. 135, 022301 (2025).
@@
- Liu et al. [2025] L. Liu, C. Zhang, J. Chen, J. Jia, X.-G. Huang, and Y.-G. Ma, Scaling approach to rigid and soft nuclear deformation through flow fluctuations in high-energy nuclear collisions (2025), arXiv:2509.09376 [nucl-th] .
@@
- Ke [2025] W. Ke, From wave-function to fireball geometry: the role of a restored broken symmetry in ultrarelativistic collisions of deformed nuclei (2025), arXiv:2509.09549 [nucl-th] .
@@
- Giacalone [2023] G. Giacalone, Many-body correlations for nuclear physics across scales: from nuclei to quark-gluon plasmas to hadron distributions, Eur. Phys. J. A 59, 297 (2023).
- Duguet et al. [2025] T. Duguet, G. Giacalone, S. Jeon, and A. Tichai, Revealing the harmonic structure of nuclear two-body correlations in high-energy heavy-ion collisions, Phys. Rev. Lett., accepted 10.1103/v2z7-wlnr (2025), arXiv:2504.02481 [nucl-th] .
- Blaizot and Giacalone [2025] J.-P. Blaizot and G. Giacalone, Angular structure of many-body correlations in atomic nuclei: From nuclear deformations to diffractive vector meson production in collisions (2025), arXiv:2504.15421 [nucl-th] .
- Sun et al. [2025] X. Sun, J. Dobaczewski, W. Nazarewicz, and H. Wibowo, Multipole tomography of composite quantum objects, (2025), to be published.
@@
- Ring and Schuck [1980] P. Ring and P. Schuck, The nuclear many-body problem (Springer-Verlag, Berlin, 1980).
- Schunck [2019] N. Schunck, ed., Energy density functional methods for atomic nuclei, IOP Expanding Physics (IOP Publishing, Bristol, UK, 2019) OCLC: 1034572493.
@@
- Zhou and Li [2025] Q. Zhou and Z. Li, Masses and quadrupole deformations of even-Z nuclei within a triaxial relativistic Hartree–Bogoliubov model, Particles 8, 57 (2025).
@@
- Feng and Wang [2024] Y. Feng and F. Wang, Review of nonflow estimation methods and uncertainties in relativistic heavy-ion collisions, J. Phys. G 52, 013001 (2024).
- Wang [2024] F. Wang, The nonflow issue in connecting anisotropy measurements to hydrodynamics in relativistic heavy-ion collisions (2024), arXiv:2402.03222 [nucl-ex] .
- STAR Collaboration [2024] STAR Collaboration, Imaging shapes of atomic nuclei in high-energy nuclear collisions, Nature 635, 67–72 (2024).
@@
- Jaiswal et al. [2025] S. Jaiswal, C. Shen, R. J. Furnstahl, U. Heinz, and M. T. Pratola, Bayesian model-data comparison incorporating theoretical uncertainties (2025), arXiv:2504.13144 [hep-ph] .